The weights of bananas in a fruit shop have a normal distribution with mean 150 grams
and standard deviation 50 grams. Three sizes of bananas are sold.
Small: under 95 grams Medium: between 95 grams and 205 grams Large: over 205 grams
(i) Find the proportion of bananas that are small. [4]
(ii) Find the weight exceeded by 10% of the bananas. [4]
The price of bananas are 10 cents, 20 cents for a medium banana and 25 cents for a
large banana.
(iii) Show that a randomly chosen banana costs 20 cents is 0.7286. [
1
Expert's answer
2020-04-01T15:37:59-0400

Let "X=" the weight of bananas in grams: "X\sim N(\mu, \sigma^2)."

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given that "\mu=150\ g,\sigma=50\ g"

(i)


"P(X<95)=P(Z<{95-150\over 50})=P(Z<-1.1)\approx0.135666"

13.5666% of bananas are small


(ii)


"P(X>X^*)=P(Z>Z^*)=0.1"


"P(Z\leq Z^*)=1-P(Z>Z^*)=1-0.1=0.9=>"

"=>Z^*\approx1.281552={X^*-150\over 50}=>X^*\approx214\ g"

The weight exceeded by 10% of the bananas is 214 grams.


(iii) The price of medium bananas are 20 cents.


"P(95<X<205)=P(X<205)-P(X<95)="

"=P(Z<{205-150\over 50})-P(Z<{95-150\over 50})="

"=P(Z<1.1)-P(Z<-1.1)\approx"

"\approx0.864334-0.135666\approx0.728668"